OpenStudy (anonymous):

Find the derivative of y with respect to the independent variable See question in comment box !

OpenStudy (anonymous):

\[y = 5^{\cos \pi \theta} \]

zepdrix (zepdrix):

Hey Rjaza! Do you remember how to take the derivative of an exponential when the base `is not e`?

zepdrix (zepdrix):

\[\Large\rm y=a^x\]\[\Large\rm y'=a^x \ln a\]This look familiar? It even applies to our e^x derivative:\[\Large\rm y=e^x\]\[\Large\rm y'=e^x \ln e\]But natural log of e is simply 1, so we don't bother writing it.

zepdrix (zepdrix):

Given this function:\[\Large\rm y=5^{stuff}\]Differentiating will give us:\[\Large\rm y'=\left(5^{stuff} \ln5\right)(stuff)'\]We have to chain rule, yes?

OpenStudy (anonymous):

@zepdrix so sorry for the delayed response!! I was working on my review :(

OpenStudy (anonymous):

my final answer is...